OpenStudy (anonymous):

Solve the equation ln(1+x^2)=1+2lnx,

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

\[\begin{align*} \ln(1+x^2)&=1+2\ln x\\ \ln(1+x^2)&=\ln e+\ln x^2\\ \ln(1+x^2)&=\ln ex^2\\ 1+x^2&=ex^2\\ &\vdots \end{align*}\]

OpenStudy (anonymous):

thank you but am not being able to get the right answer... whta answer u"ve got ?

OpenStudy (anonymous):

@iGreen help plz